Rhea, one of Saturn’s moons, has a radius of 764 π‘˜π‘š and an acceleration due to the gravity of 0.265 π‘š/𝑠 2 at its surface. Calculate its mass and average density.Β 


1
Expert's answer
2021-12-20T10:26:28-0500

Let's first find the mass of Rhea:


"g=\dfrac{GM}{R^2},""M=\dfrac{gR^2}{G}=\dfrac{0.265\ \dfrac{m}{s}\times(7.64\times10^5\ m)^2}{6.67\times10^{-11}\ \dfrac{N\times m^2}{kg^2}}=2.32\times10^{21}\ kg."

Then, we can find the volume of Rhea:


"V=\dfrac{4}{3}\pi R^3."

Finally, we can find the average density of Rhea:


"\rho=\dfrac{M}{V}=\dfrac{M}{\dfrac{4}{3}\pi R^3},""\rho=\dfrac{2.32\times10^{21}\ kg}{\dfrac{4}{3}\pi\times(7.64\times10^5\ m)^3}=1242\ \dfrac{kg}{m^3}."